OpenStudy (anonymous):

Suppose the correlation between two variables is r = 0.23. What will the new correlation be if 0.14 is added to all values of the x-variable, every value of the y-variable is doubled, and the two variables are interchanged? explain please

OpenStudy (ybarrap):

Nothing happens to r. To see this, look at the definition of correlation coefficient: $$ r_{xy}=\frac{\sum\limits_{i=1}^n (x_i-\bar{x})(y_i-\bar{y})}{(n-1) s_x s_y} =\frac{\sum\limits_{i=1}^n (x_i-\bar{x})(y_i-\bar{y})} {\sqrt{\sum\limits_{i=1}^n (x_i-\bar{x})^2 \sum\limits_{i=1}^n (y_i-\bar{y})^2}}, $$ If you add 0.14 to every term of x, both the numerator and denominator increase by this same amount and thus cancel. A similar effect happens for the other two conditions. Hope this helps.